OpenStudy (anonymous):

How do I solve 8-2b=-2/3(12b+6) Step by step explanation would be great, thank you!

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (anonymous):

@BlueMiku

OpenStudy (anonymous):

\[8-2b=-\frac{ 2 }{ 3 }(12b+6)\]

OpenStudy (crabbyoldgamer):

People aren't allowed to just give answers here.

OpenStudy (anonymous):

too complicated for me to answr, sorry. xP

OpenStudy (anonymous):

@crabbyoldgamer So do you know how to solve this equation...?

OpenStudy (smilie2015):

multiply -2/3 with (12b+6)

OpenStudy (anonymous):

so it'd come out as (-2/3 x 12b)+(-2/3x6)? can the unlike terms still multiply?

OpenStudy (smilie2015):

yes

OpenStudy (anonymous):

(12b)

OpenStudy (anonymous):

oh

OpenStudy (smilie2015):

than you will get 8-2b= -24b/3 -12/3

OpenStudy (anonymous):

taking notes hold on, you can continue if you want

OpenStudy (smilie2015):

if we simplify than -24/3 = 8 -12/3 = 4

OpenStudy (smilie2015):

so now you will get what ???

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

I have a question though

OpenStudy (smilie2015):

8-2b = -8b-4 got this step ?

OpenStudy (anonymous):

Did you only just multiply the numerator by the two whole numbers and keep the denominator the same

OpenStudy (anonymous):

?

OpenStudy (smilie2015):

look we had -2/3 (12b+6) multiply -2/3* 12b + (-2/3)*6 u can simplify in this step also... -2*4b +(-4) got it?

OpenStudy (smilie2015):